OpenStudy (calculusxy):

An oil tank contains 480 gallons of heating oil. During the month of November, 1/3 of the oil is consumed. During December 1/3 what is left is consumed. How many gallons of oil are left? @Mathis1

OpenStudy (anonymous):

Lets look at it this way. An oil tank with X gallons. In November, 1/3 is used. That means we still have 2/3 of the tank left. In December, 1/3 of what is left was consumed. So we first need to find out the fraction of the tank before we can know how many gallons are left. We know we have 2/3 left at the start of December, and used another 1/3 of that: \[\frac{ 2 }{ 3 } * \frac{ 1 }{ 3 }\] This tells us how much we used in December, but we still need to subtract the 1/3 that we used in November. \[\frac{ 2 }{ 3 }-(\frac{ 2 }{ 3 }*\frac{ 1 }{ 3 })\] This gives us the fraction of the tank left. Now we need to just times that by x number of gallons for the final answer.

OpenStudy (anonymous):

Oil consumed in November = \[\frac{1}{3} \times 480 = 160 \]gallons Remaining oil = (480 - 160 ) gallons = 320 gallons Oil consumed in November =\[ \frac{1}{3} \times 320 = 106.66 \] gallons Remaining oil after December= 320-106..66 = 213.34 gallons

OpenStudy (anonymous):

@calculusxy

OpenStudy (texaschic101):

start with 480 gallons. November, 1/3 was used so we take 1/3 of 480 = 1/3 * 480 = 480/3 = 160 480 - 160 = 320 <-- now we have 320 gallons. December, 1/3 was used 1/3 of 320 = 1/3 * 320 = 320/3 = 106.66 320 - 106.66 = 213.34 is what is left in the tank.